OpenStudy (anonymous):

Victoria is buying apples which cost $2/lb and peaches which cost $3/lb. She has $39 to spend. Let a represent the number of pounds of apples she buys, and let p represent the number of pounds of peaches she buys. Which function equation relates the number of pounds of peaches she buys to the number of pounds of apples she buys?

OpenStudy (anonymous):

A. p=-2/3a+13 B. p=-2/3a+19 1/2 C. p=-2/3a-13 D. p=-3/2a-19 1/2

OpenStudy (anonymous):

@johnweldon1993 can u help me please?

OpenStudy (anonymous):

p=-2/3a+13

OpenStudy (anonymous):

thxs:)